OpenStudy (fools101):

please help. Solve for x. 3x+6+9x=3x-14-x A) -1 B) 2 C) 4 D) -2

OpenStudy (anonymous):

12x+6=2x-14 12x-2x=-14-6 10x=-20 x=-2 D)-2

OpenStudy (fools101):

Thank u Kushashwa23!